# IFT-UAM/CSIC-24-39

### v2

@@
[…]
@@ 2 EFT framework
[…]
@@
$\text{BR}\big{(}Z\rightarrow\ell_{\alpha}^{-}\ell_{\beta}^{+}\big{)}=\dfrac{G_{F}M_{Z}^{3}}{3\sqrt{2}\pi\Gamma_{Z}}\left\{\left\lvert\frac{v^{2}}{2\Lambda^{2}}C^{\alpha\beta}_{H\ell}\right\rvert^{2}+\left\lvert\frac{v^{2}}{2\Lambda^{2}}C^{\alpha\beta}_{He}\right\rvert^{2}+\dfrac{1}{4}\left\lvert\frac{v^{2}}{2\Lambda^{2}}C^{\alpha\beta}_{eZ}\right\rvert^{2}+\dfrac{1}{4}\left\lvert\frac{v^{2}}{2\Lambda^{2}}C^{\beta\alpha}_{eZ}\right\rvert^{2}\right\},$ (19)
@@ 3 Radiative and three-body cLFV decays
BR$(\mu\to e\gamma)$ $3.1\times 10^{-13}$ MEG II (2023) [35]
BR$(\mu\to eee)$ $1.0\times 10^{-12}$ SINDRUM (1988) [36]
CR$(\mu\to e,{\rm S})$ $7.0\times 10^{-11}$ Badertscher et al. (1982) [37]
CR$(\mu\to e,{\rm Ti})$ $4.3\times 10^{-12}$ SINDRUM II (1993) [38]
CR$(\mu\to e,{\rm Pb})$ $4.6\times 10^{-11}$ SINDRUM II (1996) [39]
CR$(\mu\to e,{\rm Au})$ $7.0\times 10^{-13}$ SINDRUM II (2006) [40]
BR$(\pi^{0}\to\mu^{-}e^{+})$ $3.2\times 10^{-10}$ NA62 (2021) [41]
BR$(\pi^{0}\to\mu^{+}e^{-})$ $3.8\times 10^{-10}$ E865 (2000) [42]
BR$(\pi^{0}\to\mu e)$ $3.6\times 10^{-10}$ KTeV (2007) [43]
BR$(\eta\to\mu e)$ $6.0\times 10^{-6}$ Saturne SPES2 (1996) [44]
BR$(\eta^{\prime}\to\mu e)$ $4.7\times 10^{-4}$ CLEO (2000) [45]
BR$(\phi\to\mu e)$ $2.0\times 10^{-6}$ SND (2009) [46]
BR$(\tau\to e\gamma)$ $3.3\times 10^{-8}$ BaBar (2010) [47]
BR$(\tau\to ee\bar{e})$ $2.7\times 10^{-8}$ Belle (2010) [48]
BR$(\tau\to e\mu\bar{\mu})$ $2.7\times 10^{-8}$ Belle (2010) [48]
BR$(\tau\to e\pi)$ $8.0\times 10^{-8}$ Belle (2007) [49]
BR$(\tau\to e\eta)$ $9.2\times 10^{-8}$ Belle (2007) [49]
BR$(\tau\to e\eta^{\prime})$ $1.6\times 10^{-7}$ Belle (2007) [49]
BR$(\tau\to e\pi\pi)$ $2.3\times 10^{-8}$ Belle (2012) [50]
BR$(\tau\to e\omega)$ $2.4\times 10^{-8}$ Belle (2023) [51]
BR$(\tau\to e\phi)$ $2.0\times 10^{-8}$ Belle (2023) [51]
BR$(\tau\to\mu\gamma)$ $4.2\times 10^{-8}$ Belle (2021) [52]
BR$(\tau\to\mu\mu\bar{\mu})$ $2.1\times 10^{-8}$ Belle (2010) [48]
BR$(\tau\to\mu e\bar{e})$ $1.8\times 10^{-8}$ Belle (2010) [48]
BR$(\tau\to\mu\pi)$ $1.1\times 10^{-7}$ BaBar (2006) [53]
BR$(\tau\to\mu\eta)$ $6.5\times 10^{-8}$ Belle (2007) [49]
BR$(\tau\to\mu\eta^{\prime})$ $1.3\times 10^{-7}$ Belle (2007) [49]
BR$(\tau\to\mu\pi\pi)$ $2.1\times 10^{-8}$ Belle (2012) [50]
BR$(\tau\to\mu\omega)$ $3.9\times 10^{-8}$ Belle (2023) [51]
BR$(\tau\to\mu\phi)$ $2.3\times 10^{-8}$ Belle (2023) [51]
BR$(\tau\to ee\bar{\mu})$ $1.5\times 10^{-8}$ Belle (2010) [48]
BR$(\tau\to\mu\mu\bar{e})$ $1.7\times 10^{-8}$ Belle (2010) [48]
@@ 4 Semileptonic cLFV observables
$\text{BR}\left(\tau\rightarrow\ell\pi^{0}\right)=\dfrac{G_{F}^{2}f_{\pi}^{2}(m_{\tau}^{2}-m_{\pi}^{2})^{2}}{8\pi\Gamma_{\tau}m_{\tau}}\sum_{X=L,R}\left\lvert c^{uA}_{\tau\ell X}-c^{dA}_{\tau\ell X}+\dfrac{m_{\pi}^{2}}{m_{\tau}\left(m_{u}+m_{d}\right)}\left(c_{\tau\ell X}^{uP}-c_{\tau\ell X}^{dP}\right)\right\rvert^{2},$ (24)
@@
[…]
@@ 4.2 $\boldsymbol{\mu-e}$ conversion in nuclei
[…]
@@
where the fit functions ±$\mathcal{F}^{\Sigma^{{}^{\prime}(^{\prime\prime})}_{1}}_{\pm}(u)$ are given in Ref. [29].
Consequently, for each lepton chirality and regardless on how many isotopes we have data on, at most 4 directions can be constrained by SD conversion, namely the the 4 WC combinations in Eqs. (35)-(36): $C^{0\mathcal{T}}_{X}$, $C^{1\mathcal{T}}_{X}$, $C^{0\mathcal{L}}_{X}$ and $C^{1\mathcal{L}}_{X}$. Consequently, the parameter space of axial, pseudoscalar and tensor operators cannot be fully probed with only SD $\mu-e$ conversion and complementary probes, such as meson decays, are needed [^8] This seems at odds with the conclusions of Ref. [29]. However, even though higher multipoles were considered in that work, the only combinations of operators contributing are the four in Eqs. (35)-(36). Thus, even these higher orders are not able to provide additional information..
@@ 4.3 Meson decays: $\boldsymbol{M\rightarrow\mu e}$
$\text{BR}\left(\pi^{0}\rightarrow\mu^{\mp}e^{\pm}\right)=\dfrac{G_{F}^{2}f_{\pi}^{2}m_{\mu}^{2}\big{(}m_{\pi}^{2}-m_{\mu}^{2}\big{)}^{2}}{4\pi\Gamma_{\pi^{0}}m_{\pi}^{3}}$ $\sum_{X=L,R}\Bigg{|}c^{uA}_{\mu eX}-c^{dA}_{\mu eX}\mp\frac{m_{\pi}^{2}}{2m_{\mu}}\left(\frac{c^{uP}_{\mu eX}}{m_{u}}-\frac{c^{dP}_{\mu eX}}{m_{d}}\right)\Bigg{|}^{2},$ (43)
$\text{BR}\left(\eta\rightarrow\mu^{\mp}e^{\pm}\right)=\dfrac{G_{F}^{2}f_{\pi}^{2}m_{\mu}^{2}\big{(}m_{\eta}^{2}-m_{\mu}^{2}\big{)}^{2}}{4\pi\Gamma_{\eta}m_{\eta}^{3}}$ $\sum_{X=L,R}\Bigg{|}\dfrac{f_{\eta}^{u}}{f_{\pi}}(c^{uA}_{\mu eX}+c^{dA}_{\mu eX})+\dfrac{f_{\eta}^{s}}{f_{\pi}}c^{sA}_{\mu eX}$
@@
$\text{BR}\left(\eta^{\prime}\rightarrow\mu^{\mp}e^{\pm}\right)=\dfrac{G_{F}^{2}f_{\pi}^{2}m_{\mu}^{2}\big{(}m_{\eta^{\prime}}^{2}-m_{\mu}^{2}\big{)}^{2}}{4\pi\Gamma_{\eta^{\prime}}m_{\eta^{\prime}}^{3}}$ $\sum_{X=L,R}\Bigg{|}\dfrac{f_{\eta^{\prime}}^{u}}{f_{\pi}}(c^{uA}_{\mu eX}+c^{dA}_{\mu eX})+\dfrac{f_{\eta^{\prime}}^{s}}{f_{\pi}}c^{sA}_{\mu eX}$
@@ 5.2 $\boldsymbol{\mu-e}$ sector
In complete analogy to the $\tau-\ell$ sector, there are $15$ independent LEFT operators involving $\mu$, $e$ and two light quarks. In an idealized scenario, i.e. neglecting nuclear uncertainties, $\mu-e$ conversion data can probe $8$ different combinations of WCs, as outlined in section 4.2, 4 from SI and 4 from SD $\mu-e$ conversion. Additionally, the $\pi^{0}$ and $\eta$ decays discussed in section 4.3 can provide (weaker) bounds on 2 additional directions [^10] The decays to $\mu^{-}e^{+}$ and $\mu^{+}e^{-}$ would in principle lead to 4 additional constraints. However, the only meaningful constraint derived from the $\eta$ decay is on the pseudoscalar combination of operators, as the bound on the axial contribution is too weak given the chirality suppression and one of the bounds coming from the $\pi^{0}$ decay is redundant with the constraints from SD $\mu-e$ conversion.. Thus, $15-10=5$ flat directions are present in this scenario. Nevertheless, as for the $\tau$ case, some Lorentz combinations can be fully constrained even in presence of these flat directions.
@@ 6 SMEFT analysis
$c^{sP}_{\alpha\beta X}$ =±$=\pm c^{sS}_{\alpha\beta X}\,,$ (49)
@@ 6.2 $\boldsymbol{\mu-e}$ sector
    Nuclear overlap integrals: these define the directions probed by SI $\mu-e$ conversion. The main source of uncertainty comes from the fact that lepton-nucleon interactions are computed at LO in $\chi$PT, see Refs. [18, 82, 66] for discussions about the size of the possible NLO corrections. We will consider $5\%$ and $10\%$ uncertainties for the overlap integrals of light and heavy nuclei, respectively. In particular, we will consider the parameters $V^{(p),(n)}$ and $S^{(p),(n)}$ in Eq. (31) as free and independent with Gaussian priors centered around their nominal value and a $5\%$/$10\%$ uncertainty. With this parametrization, there will be a confidence level at which the freedom allowed for the overlap integrals makes two of them parallel [80] and, therefore, redundant instead of complementary at this C.L. This has a dramatic impact in the analysis, since a new flat direction appears, loosening many constraints and changing significantly the correlations among the remaining ones, as depicted in Appendix A.
@@ 7.2 $\boldsymbol{\mu-e}$ sector
The results of our fit taking into account nuclear uncertainties, are shown in Fig. 5 as dark red bars. The treatment of the nuclear uncertainties as nuisance parameters is equivalent to what is described in section 6.2. We again find that bounds on vector and scalar operators are the most affected, as they can no longer be solely constrained through SI $\mu-e$ conversion. However, one important difference is that now the constraints have suffered from a milder degradation with respect to the case of Fig. 4. This is a consequence of the fact that, in this simplified scenario, $\mu-e$ conversion can place constraints on $8-2=6$ combinations after nuclear uncertainties are taken into account. Therefore, in order to completely constrain the $7$ WCs, only one additional constraint coming from meson decays is necessary, contrary to the scenario with $s$-quarks, in which two extra meson constraints were necessary. As in the previous section, these constraints are very strongly correlated and are significantly different between the 1 and 2 $\sigma$ regions, when the nuisance parameters are able to reduce the number of constrained directions, as show in Fig. 8 in appendix A.
@@ Acknowledgments.
The authors thank Javier Menéndez, Jacobo López-Pavón and Salvador Urrea for very illuminating discussions. This project has received support from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 860881-HIDDeN and No 101086085 - ASYMMETRY, and from the Spanish Research Agency (Agencia Estatal de Investigación) through the Grant IFT Centro de Excelencia Severo Ochoa No CEX2020-001007-S and Grant PID2022 137127NB-I00 funded by MCIN/AEI/10.13039/501100011033. We acknowledge support from the HPC-Hydra cluster at IFT. XM acknowledges funding from the European Union’s Horizon Europe Programme under the Marie Skłodowska-Curie grant agreement no. 101066105-PheNUmenal. The work of DNT was supported by the Spanish MIU through the National Program FPU (grant number FPU20/05333).
@@ Appendix B Bounds on dipole and four-lepton operators
$\begin{pmatrix}c_{e\mu}^{e\gamma}\\ \\ c_{e\tau}^{e\gamma}\\ \\ c_{\mu\tau}^{e\gamma}\end{pmatrix}<\begin{pmatrix}6.0\times 10^{-12}\\ \\ 7.9\times 10^{-8}\\ \\ 8.9\times 10^{-8}\end{pmatrix}\,,\hskip 42.67912pt\begin{pmatrix}c_{e\mu L}^{eeLV}\\ \\ c_{e\tau L}^{eeLV}\\ \\ c_{\mu\tau L}^{\mu\mu LV}\\ \\ c_{e\tau L}^{\mu\mu LV}\\ \\ c_{\mu\tau L}^{eeLV}\\ \\ c_{e\tau L}^{e\mu LV}\\ \\ c_{\mu\tau L}^{\mu eLV}\end{pmatrix}<\begin{pmatrix}6.2\times 10^{-6}\\ \\ 2.4\times 10^{-3}\\ \\ 2.1\times 10^{-3}\\ \\ 2.0\times 10^{-3}\\ \\ 2.0\times 10^{-3}\\ \\ 1.8\times 10^{-3}\\ \\ 1.9\times 10^{-3}\end{pmatrix}\,,$ (61)
$\begin{pmatrix}c_{e\mu L}^{eeRV}\\ \\ c_{e\tau L}^{eeRV}\\ \\ c_{\mu\tau L}^{\mu\mu RV}\\ \\ c_{e\tau L}^{\mu\mu RV}\\ \\ c_{\mu\tau L}^{e\mu RV}\\ \\ c_{\mu\tau L}^{eeRV}\\ \\ c_{e\tau L}^{\mu eRV}\\ \\ c_{e\tau L}^{e\mu RV}\\ \\ c_{\mu\tau L}^{\mu eRV}\end{pmatrix}<\begin{pmatrix}5.2\times 10^{-6}\\ \\ 2.0\times 10^{-3}\\ \\ 1.8\times 10^{-3}\\ \\ 2.0\times 10^{-3}\\ \\ 2.0\times 10^{-3}\\ \\ 2.0\times 10^{-3}\\ \\ 2.0\times 10^{-3}\\ \\ 1.5\times 10^{-3}\\ \\ 1.6\times 10^{-3}\end{pmatrix}\,,\hskip 42.67912pt\begin{pmatrix}c_{e\mu R}^{eeRS}\\ \\ c_{e\tau R}^{eeRS}\\ \\ c_{\mu\tau R}^{\mu\mu RS}\\ \\ c_{e\tau R}^{\mu\mu RS}\\ \\ c_{\mu\tau R}^{e\mu RS}\\ \\ c_{\mu\tau R}^{eeRS}\\ \\ c_{e\tau R}^{\mu eRS}\\ \\ c_{e\tau R}^{e\mu RS}\\ \\ c_{\mu\tau R}^{\mu eRS}\end{pmatrix}<\begin{pmatrix}3.1\times 10^{-6}\\ \\ 1.2\times 10^{-3}\\ \\ 1.1\times 10^{-3}\\ \\ 1.4\times 10^{-3}\\ \\ 1.4\times 10^{-3}\\ \\ 1.4\times 10^{-3}\\ \\ 1.4\times 10^{-3}\\ \\ 9.0\times 10^{-4}\\ \\ 9.6\times 10^{-4}\end{pmatrix}\,.$ (62)

### v3

@@
Enrique Fernández-Martínez***enrique.fernandez@csic.es, Xabier Marcanoxabier.marcano@uam.es and Daniel Naredo-Tuerodaniel.naredo@uam.es
@@ 2 EFT framework
Table 1: Dimension-66 SMEFT operators relevant for cLFV processes, with flavour indices $\alpha,\beta,\gamma,\delta$. $Q$ and $L$ are quark and lepton $SU(2)_{L}$ doublets ($a,b=1,2$, $I=1,2,3$ $SU(2)_{L}$ indices and $\tau^{I}$ the Pauli matrices), while $u$, $d$ and $e$ are up, down quark and lepton singlets. $H$ denotes the Higgs doublet with $\mbox{${H^{\dagger}\,\raisebox{4.2679pt}{${}^{\leftrightarrow}$}\hskip-11.38109ptD_{\mu}\,H}$}\equiv H^{\dagger}(D_{\mu}H)-(D_{\mu}H)^{\dagger}H$, and $B_{\mu\nu}$ and $W^{I}_{\mu\nu}$ are the $U(1)_{Y}$ and $SU(2)_{L}$ field strengths, respectively.
@@
$\hskip-9.10509pt\text{BR}\big(Z\rightarrow\ell_{\alpha}^{-}\ell_{\beta}^{+}\big)=\dfrac{G_{F}M_{Z}^{3}}{3\sqrt{2}\pi\Gamma_{Z}}\hskip-2.84544pt\left\{\left\lvert\frac{v^{2}}{2\Lambda^{2}}C^{\alpha\beta}_{H\ell}\right\rvert^{2}+\left\lvert\frac{v^{2}}{2\Lambda^{2}}C^{\alpha\beta}_{He}\right\rvert^{2}+\dfrac{1}{4}\left\lvert\frac{v^{2}}{2\Lambda^{2}}C^{\alpha\beta}_{eZ}\right\rvert^{2}+\dfrac{1}{4}\left\lvert\frac{v^{2}}{2\Lambda^{2}}C^{\beta\alpha}_{eZ}\right\rvert^{2}\right\},$ (19)
@@ 3 Radiative and three-body cLFV decays
[…]
@@ 4 Semileptonic cLFV observables
$\hskip-6.82881pt\text{BR}\left(\tau\rightarrow\ell\pi^{0}\right)=\dfrac{G_{F}^{2}f_{\pi}^{2}(m_{\tau}^{2}-m_{\pi}^{2})^{2}}{8\pi\Gamma_{\tau}m_{\tau}}\hskip-7.11317pt\sum_{X=L,R}\left\lvert c^{uA}_{\tau\ell X}-c^{dA}_{\tau\ell X}+\dfrac{m_{\pi}^{2}}{m_{\tau}\left(m_{u}+m_{d}\right)}\left(c_{\tau\ell X}^{uP}-c_{\tau\ell X}^{dP}\right)\right\rvert^{2},$ (24)
@@
$\text{BR}\left(\tau\rightarrow\ell\phi\right)=\dfrac{G_{F}^{2}f_{\phi}^{2}m_{\phi}^{2}(m_{\tau}^{2}-m_{\phi}^{2})}{4\pi\Gamma_{\tau}m_{\tau}}\sum_{X=L,R}\Bigg\{$ $\left(\frac{m_{\tau}^{2}}{m_{\phi}^{2}}+1-2\frac{m_{\phi}^{2}}{m_{\tau}^{2}}\right)\left\lvert c^{sV}_{\tau\ell X}\right\rvert^{2}$
@@ 4.2 $\boldsymbol{\mu-e}$ conversion in nuclei
$c^{0x}_{X}=\dfrac{c^{px}_{X}+c^{nx}_{X}}{2}\,,\hskip 28.45274ptc^{1x}_{X}=\dfrac{c^{px}_{X}-c^{nx}_{X}}{2}\,,\hskip 14.22636pt\text{for}\hskip 14.22636ptX=L,R,\hskip 14.22636ptx=A,P,T,$ (37)
@@
where the fit functions $\mathcal{F}^{\Sigma^{{}^{\prime}(^{\prime\prime})}_{1}}_{\pm}(u)$ are given in Ref. [29].
Consequently, for each lepton chirality and regardless on how many isotopes we have data on, at most 4 directions can be constrained by SD conversion, namely the 4 WC combinations in Eqs. (35)-(36): $C^{0\mathcal{T}}_{X}$, $C^{1\mathcal{T}}_{X}$, $C^{0\mathcal{L}}_{X}$ and $C^{1\mathcal{L}}_{X}$. Consequently, the parameter space of axial, pseudoscalar and tensor operators cannot be fully probed with only SD $\mu-e$ conversion and complementary probes, such as meson decays, are needed [^8] This seems at odds with the conclusions of Ref. [29]. However, even though higher multipoles were considered in that work, the only combinations of operators contributing are the four in Eqs. (35)-(36). Thus, even these higher orders are not able to provide additional information..
@@ 4.3 Meson decays: $\boldsymbol{M\rightarrow\mu e}$
$\text{BR}\left(\pi^{0}\rightarrow\mu^{\mp}e^{\pm}\right)=\dfrac{G_{F}^{2}f_{\pi}^{2}m_{\mu}^{2}\big(m_{\pi}^{2}-m_{\mu}^{2}\big)^{2}}{4\pi\Gamma_{\pi^{0}}m_{\pi}^{3}}\hskip-28.45274pt$ $\sum_{X=L,R}\Bigg|c^{uA}_{\mu eX}-c^{dA}_{\mu eX}\mp\frac{m_{\pi}^{2}}{2m_{\mu}}\left(\frac{c^{uP}_{\mu eX}}{m_{u}}-\frac{c^{dP}_{\mu eX}}{m_{d}}\right)\Bigg|^{2},$ (43)
$\text{BR}\left(\eta\rightarrow\mu^{\mp}e^{\pm}\right)=\dfrac{G_{F}^{2}f_{\pi}^{2}m_{\mu}^{2}\big(m_{\eta}^{2}-m_{\mu}^{2}\big)^{2}}{4\pi\Gamma_{\eta}m_{\eta}^{3}}\hskip-28.45274pt$ $\sum_{X=L,R}\Bigg|\dfrac{f_{\eta}^{u}}{f_{\pi}}(c^{uA}_{\mu eX}+c^{dA}_{\mu eX})+\dfrac{f_{\eta}^{s}}{f_{\pi}}c^{sA}_{\mu eX}$
@@
$\text{BR}\left(\eta^{\prime}\rightarrow\mu^{\mp}e^{\pm}\right)=\dfrac{G_{F}^{2}f_{\pi}^{2}m_{\mu}^{2}\big(m_{\eta^{\prime}}^{2}-m_{\mu}^{2}\big)^{2}}{4\pi\Gamma_{\eta^{\prime}}m_{\eta^{\prime}}^{3}}\hskip-21.33955pt$ $\sum_{X=L,R}\Bigg|\dfrac{f_{\eta^{\prime}}^{u}}{f_{\pi}}(c^{uA}_{\mu eX}+c^{dA}_{\mu eX})+\dfrac{f_{\eta^{\prime}}^{s}}{f_{\pi}}c^{sA}_{\mu eX}$
@@ 5.2 $\boldsymbol{\mu-e}$ sector
In complete analogy to the $\tau$-$\ell$ sector, there are $15$ independent LEFT operators involving $\mu$, $e$ and two light quarks. In an idealized scenario, i.e. neglecting nuclear uncertainties, $\mu-e$ conversion data can probe $8$ different combinations of WCs, as outlined in section 4.2, 4 from SI and 4 from SD $\mu-e$ conversion. Additionally, the $\pi^{0}$ and $\eta$ decays discussed in section 4.3 can provide (weaker) bounds on 2 additional directions [^10] The decays to $\mu^{-}e^{+}$ and $\mu^{+}e^{-}$ would in principle lead to 4 additional constraints. However, the only meaningful constraint derived from the $\eta$ decay is on the pseudoscalar combination of operators, as the bound on the axial contribution is too weak given the chirality suppression and one of the bounds coming from the $\pi^{0}$ decay is redundant with the constraints from SD $\mu-e$ conversion.. Thus, 1510=515-10=5 flat directions are present in this scenario. Nevertheless, as for the $\tau$ case, some Lorentz combinations can be fully constrained even in presence of these flat directions.
@@ 6 SMEFT analysis
$c^{sP}_{\alpha\beta X}$ $=\pm c^{sS}_{\alpha\beta X}\,$, (49)
@@ 6.2 $\boldsymbol{\mu-e}$ sector
[…]
@@ 7.2 $\boldsymbol{\mu-e}$ sector
The results of our fit taking into account nuclear uncertainties, are shown in Fig. 5 as dark red bars. The treatment of the nuclear uncertainties as nuisance parameters is equivalent to what is described in section 6.2. We again find that bounds on vector and scalar operators are the most affected, as they can no longer be solely constrained through SI $\mu-e$ conversion. However, one important difference is that now the constraints have suffered from a milder degradation with respect to the case of Fig. 4. This is a consequence of the fact that, in this simplified scenario, $\mu-e$ conversion can place constraints on 82=68-2=6 combinations after nuclear uncertainties are taken into account. Therefore, in order to completely constrain the $7$ WCs, only one additional constraint coming from meson decays is necessary, contrary to the scenario with $s$-quarks, in which two extra meson constraints were necessary. As in the previous section, these constraints are very strongly correlated and are significantly different between the 1 and 2 $\sigma$ regions, when the nuisance parameters are able to reduce the number of constrained directions, as show in Fig. 8 in appendix A.
@@ Acknowledgments.
The authors thank Javier Menéndez, Jacobo López-Pavón and Salvador Urrea for very illuminating discussions, and Xavier Ponce Díaz for pointing out a typo in our numerical analysis of the three-body leptonic decays. This project has received support from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 860881-HIDDeN and No 101086085 - ASYMMETRY, and from the Spanish Research Agency (Agencia Estatal de Investigación) through the Grant IFT Centro de Excelencia Severo Ochoa No CEX2020-001007-S and Grant PID2022 137127NB-I00 funded by MCIN/AEI/10.13039/501100011033. We acknowledge support from the HPC-Hydra cluster at IFT. XM acknowledges funding from the European Union’s Horizon Europe Programme under the Marie Skłodowska-Curie grant agreement no. 101066105-PheNUmenal. The work of DNT was supported by the Spanish MIU through the National Program FPU (grant number FPU20/05333).
@@ Appendix B Bounds on dipole and four-lepton operators
$\begin{pmatrix}c_{e\mu}^{e\gamma}\\ \\ c_{e\tau}^{e\gamma}\\ \\ c_{\mu\tau}^{e\gamma}\end{pmatrix}<\begin{pmatrix}6.0\times 10^{-12}\\ \\ 7.9\times 10^{-8}\\ \\ 8.9\times 10^{-8}\end{pmatrix}\,,\hskip 42.67912pt\begin{pmatrix}c_{e\mu L}^{eeLV}\\ \\ c_{e\tau L}^{eeLV}\\ \\ c_{\mu\tau L}^{\mu\mu LV}\\ \\ c_{e\tau L}^{\mu\mu LV}\\ \\ c_{\mu\tau L}^{eeLV}\\ \\ c_{e\tau L}^{e\mu LV}\\ \\ c_{\mu\tau L}^{\mu eLV}\end{pmatrix}<\begin{pmatrix}7.8\times 10^{-7}\\ \\ 3.0\times 10^{-4}\\ \\ 2.7\times 10^{-4}\\ \\ 4.3\times 10^{-4}\\ \\ 4.3\times 10^{-4}\\ \\ 2.3\times 10^{-4}\\ \\ 2.4\times 10^{-4}\end{pmatrix}\,$, (61)
$\begin{pmatrix}c_{e\mu L}^{eeRV}\\ \\ c_{e\tau L}^{eeRV}\\ \\ c_{\mu\tau L}^{\mu\mu RV}\\ \\ c_{e\tau L}^{\mu\mu RV}\\ \\ c_{\mu\tau L}^{e\mu RV}\\ \\ c_{\mu\tau L}^{eeRV}\\ \\ c_{e\tau L}^{\mu eRV}\\ \\ c_{e\tau L}^{e\mu RV}\\ \\ c_{\mu\tau L}^{\mu eRV}\end{pmatrix}<\begin{pmatrix}1.1\times 10^{-6}\\ \\ 4.3\times 10^{-4}\\ \\ 3.8\times 10^{-4}\\ \\ 4.3\times 10^{-4}\\ \\ 4.3\times 10^{-4}\\ \\ 4.3\times 10^{-4}\\ \\ 4.3\times 10^{-4}\\ \\ 3.2\times 10^{-4}\\ \\ 3.4\times 10^{-4}\end{pmatrix}\,,\hskip 42.67912pt\begin{pmatrix}c_{e\mu R}^{eeRS}\\ \\ c_{e\tau R}^{eeRS}\\ \\ c_{\mu\tau R}^{\mu\mu RS}\\ \\ c_{e\tau R}^{\mu\mu RS}\\ \\ c_{\mu\tau R}^{e\mu RS}\\ \\ c_{\mu\tau R}^{eeRS}\\ \\ c_{e\tau R}^{\mu eRS}\\ \\ c_{e\tau R}^{e\mu RS}\\ \\ c_{\mu\tau R}^{\mu eRS}\end{pmatrix}<\begin{pmatrix}3.1\times 10^{-6}\\ \\ 1.2\times 10^{-3}\\ \\ 1.1\times 10^{-3}\\ \\ 8.6\times 10^{-4}\\ \\ 8.6\times 10^{-4}\\ \\ 8.6\times 10^{-4}\\ \\ 8.6\times 10^{-4}\\ \\ 9.0\times 10^{-4}\\ \\ 9.6\times 10^{-4}\end{pmatrix}\,$. (62)